OpenStudy (anonymous):

What is the discontinuity of the function f(x) = the quantity of x squared minus 3 x minus 28, all over x plus 4

OpenStudy (mathmale):

Here's my interpretation of what you've presented:\[\frac{ x ^{2}-3x-28 }{ x+4 }=\frac{ (x-7)(x+4) }{ x+4 }\]

OpenStudy (mathmale):

Although you can cancel the 2 (x+4) terms, that does not remove the discontinuity you'll have at x=-4. If you have any questions about this, by all means please ask.

OpenStudy (anonymous):

its not right

OpenStudy (anonymous):

its f(x)= then u can read it

OpenStudy (mathmale):

Explain what you mean, please, by "its not right."

OpenStudy (mathmale):

While it's better to label a function like this with f(x)=, as you suggest, that will have no bearing on the answer to this question.

OpenStudy (anonymous):

x^2-3x-28 all over x +4

OpenStudy (anonymous):

sorry about the bad communication

OpenStudy (anonymous):

wait you are right

OpenStudy (anonymous):

but how do you find the discontinuity